OpenStudy (anonymous):

What is the equation of a line that passes through (7, 8) and has a slope of -3?

OpenStudy (ccswims):

first you must do point-slope form before we can do the equation

OpenStudy (ccswims):

y - y1 = m(x - x1) m = slope and we know -3 is our slope

OpenStudy (ccswims):



OpenStudy (ccswims):

can you distribute the -3 to x and 7 by yourself for now?

OpenStudy (anonymous):

Thanks soo much